# On the Brun–Titchmarsh theorem

### v1

@@ Abstract.
The classical Brun–Titchmarsh theorem gives an upper bound, which is of correct order of magnitude in the full range, for the number of primes $p\leqslant x$ satisfying $p\equiv a\,(\operatorname{mod}{q})$. We strength this inequality for $\log q/\log x$ in different ranges, improving upon previous works by Motohashi, Goldfeld, Iwaniec, Friedlander and Iwaniec, and Maynard for general or special moduli. In particular, we obtain a Burgess-like constant in the full range $q<x^{1/2-}.$ The proof is based on various estimates for character and exponential sums, by appealing to arithmetic exponent pairs and bilinear forms with algebraic trace functions from $\ell$-adic cohomology, and sums of Kloosterman sums from spectral theory of automorphic forms, as well as large value theorem for Dirichlet polynomials due to Huxley.
@@ 1.1. Main results
In this paper, we establish several choices of $C(\varpi)$ for general moduli and smooth moduli $q\sim x^{\varpi}$. We now quickly state our results.
@@
Denote by $\vartheta$ the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$ Then we may take
(1.4) $C(\varpi)=\frac{16}{8-(3+2\vartheta)\varpi}$
@@
[…]
@@
holds for any non-trivial Dirichlet character modulo $q$ and $r\in\{1,2,3\}$. Taking $r=2$, (1.5) implies the so-called Burgess bound $L(\frac{1}{2},\chi)\ll q^{3/16+\varepsilon}$ for Dirichlet $L$-functions. The restriction $\varpi<9/20$ is quite essential in Iwaniec’s approach, and Theorem 1.1 modulo the Ramanujan–Petersson conjecture extends the validity of Iwaniec’s constant in a much wider range. Different from [Iw82], our approach relies heavily on the well-factorization of the linear sieve weight constructed by Iwaniec [Iw80] and Kloostermania developed by Deshouillers and Iwaniec [DI82]. In the prime moduli case, the above choice for $C(\varpi)$ is independent from the Ramanujan–Petersson conjecture, and this will follow from another different approach, by appealing to a recent work of Blomer, Humphries, Khan and Milinovich [BHKM20] on twisted fourth moment of Dirichlet $L$-functions to prime moduli. To echo the above convention of Burgess-like constant, we should mention that the bound in [BHKM20] implies the above Burgess bound for $L(\frac{1}{2},\chi)$ from an approach completely different from character sums; see Proposition 1.1 below.
Note that Iwaniec [Iw82] is in fact able to take $C(\varpi)=6/(3-\varpi)$ for $\varpi<9/20,$ which corresponds to $r=3$ in (1.5).
@@
$C(\varpi)=\begin{cases}8/(5-5\varpi),\quad&\varpi\in[1/2,~{}12/23[,\\ 32/(32-43\varpi),\quad&\varpi\in[12/23,~{}32/61[,\\ 24/(16-17\varpi),&\varpi\in[32/61,~{}8/15[,\\ 48/(40-49\varpi),&\varpi\in[8/15,~{}7/13[,\\ 16/(11-12\varpi),&\varpi\in[7/13,~{}6/11[,\\ 32/(28-35\varpi),&\varpi\in[6/11,~{}4/7[.\end{cases}$
Remark 1.
We may relax $q$ to be squarefree by using the work of Löffel [Lo16], as an extension of Lemma 4.3. It is also possible to generalize Theorem 1.2 to arbitrary moduli along with the same line.
@@
]9/20,2/3]]9/20,2/3]] 9 / 20 , 2 / 3 ] $8/(6-7\varpi)$ Iwaniec [Iw82, Theorem 6]
@@
The second part of this paper is to reduce the admissible value of $C(\varpi)$ further by assuming that $q$ has very good factorizations. For a small constant $\eta>0$, we say $q$ is $\eta$-smooth if all prime factors of $q$ do not exceed $q^{\eta}.$ The assumption on the smoothness allows us to utilize the factorization of $q$ to control the resultant algebraic exponential sums very effectively. In particular, one may go far beyond the so-called Pólya–Vinogradov barrier for short exponential and character sums, as a consequence of the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs developed by Wu and Xi [WX21]. One is referred to Section 4.2 for details.
@@
$C(\varpi)=\frac{160}{89-91\varpi}$
@@
$\frac{9}{51}\leqslant\varpi\leqslant\frac{9}{11}=\frac{2}{3}+\frac{5}{33},$
provided that $\eta$ is small enough.
Theorem 1.3 is in fact a special case of the following slightly general formulation on taking $(\kappa,\lambda)=(\tfrac{1}{20},\tfrac{33}{40})$.
@@
$f(\kappa,\lambda)$ $=\frac{1+\kappa-\lambda}{1+2\kappa-\lambda},$
$g_{\varpi}(\kappa,\lambda)$ $=(3+\kappa-\lambda)-(3+2\kappa-\lambda)\varpi.$
@@
in which case we get $g_{\varpi}(\kappa,\lambda)=0.723659\cdots$. This yields $C(2/3)=4/g_{\varpi}(\kappa,\lambda)\approx 5.2746$ in Theorem 1.4. Note that van Lint and Richert [LR65] obtained the constant $62/(1-\varpi)=$ 6 for $\varpi=2/3.$
The following theorem turns out to be better than Theorem 1.5 when $\varpi$ is a bit smaller than $1/2$.
@@
Theorem 1.5 can be compared with the following conditional result due to Motohashi [Mo74]. Under the Linderlöf Hypothesis for Dirichlet $L$-functions, he proved that take
@@
The following two figures present our new choices of $C(\varpi)$ for $\varpi<1/2$ and $\varpi>1/2,$ respectively, and one may see when we can succeed or not.
@@
1.2. Conditional results
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}|L(\tfrac{1}{2}+it,\chi)|^{4}\ll q^{\varepsilon}(1+|t|)^{O(1)}$
with $\delta>0$ as large as possible. This motivates us to make the following conjecture.
Conjecture 1.1.
[…]
By extending Motohashi’s identity to link fourth moment of Dirichlet $L$-functions and cubic moment of $\mathrm{GL}_{2}$ $L$-functions, Blomer, Humphries, Khan and Milinovich [BHKM20] proved the following bound.
Proposition 1.1.
[…]
The following two theorems show that Conjecture 1.1 may guarantee sharp choices of $C(\varpi)$ for $\varpi<1/2$, by combining the possible good factorizations of $q$.
Theorem 1.6.
[…]
$C(\varpi)=\begin{cases}2,\quad&\varpi\in[1/8,~{}10/(24-5\delta)[,\\ 20/(20-(24-5\delta)\varpi),&\varpi\in[10/(24-5\delta),~{}9/20[,\end{cases}$
provided that $\eta$ is small enough.
Remark 2.
Theorem 1.5 corresponds to the degenerate case $\delta=0$ of Theorem 1.6.
Theorem 1.7.
Assume that Conjecture 1.1 holds for all $q\in\mathbf{Z}^{+}$ and some $\delta\in[0,16/45[.$ Then we may take
$C(\varpi)=\frac{4}{2-(1-\delta)\varpi}$
for $\varpi\in[9/20,1/2].$
@@ 1.3. Ingredients of proof and comparisons with previous works
  • Both of van Lint and Richert [LR65] and Motohashi [Mo74, Mo73] utilized Selberg’s sieve, which is more convenient than Brun’s sieve at that time. Recall that one takes $C(\varpi)=2/(1-\varpi)$ in [LR65], and the constant $2$ is resulted naturally by applications of the linear sieve with its best possible main term. On the other hand, Montgomery and Vaughan [MV73] developed the large sieve, from which they deduced that $\pi(x;q,a)<2x/(\varphi(q)\log(q/x))$ for all $1\leqslant q<x$ and $(a,q)=1$. The above denominator $(1-\varpi)$ comes from the level $(x/q)^{1-\varepsilon}$ after a trivial treatment of error terms. It was Motohashi [Mo74, Mo73] who first dealt with the error term in a non-trivial manner, and he successfully employed the bilinear structure in Selberg’s sieve. He employed multiplicative characters, and a basic tool is the mean value of Dirichlet $L$-functions, which prevents him from improving van Lint and Richert [LR65] for $\varpi>1/2.$
  • Iwaniec [Iw82] developed a new linear sieve of combinatorial type with well-factorable weights. This far-reaching work was originally motivated by Chen [Ch75] and Motohashi [Mo74], and the merit is that one may optimize the balances between convolutions in the error term, and the sign changes among them can be captured more effectively. In particular, this well-factorization in the error term allows Iwaniec [Iw82] to employ (Weil’s bound for) Kloosterman sums to beat the barrier $\varpi\approx 1/2$. He was able to beat van Lint and Richert [LR65] even if $\varpi$ is close to $2/3-.$
@@
Note that $(c_{r})$ has a convolution structure, coming from Iwaniec’s new construction of the Rosser–Iwaniec sieve. We utilize the special structure of $(c_{r})$ as efficiently as possible, and for $k=2$ we may give a another kind of Linderlöf Hypothesis on average for $R\approx q^{(5-2\vartheta)/8}$. Together with large sieves (Lemma 3.1) and large value estimate for Dirichlet polynomials (Lemma 3.2), this can also produce a level of distribution in linear sieves with the same strength as the above approach. The advantage is that $q$ can be an arbitrary large integer, but we are restricted by the possibility of exceptional spectrum of the non-Eucliean Laplacian. Modulo the Ramanujan–Petersson conjecture, we reach exactly the same level.
@@
(1) If $q$ is squarefree and has only very small prime factors, the factorization of $q$ can be controlled very well, so that the work on arithmetic exponent pairs, developed by Wu and Xi [WX21], can be applied. In such situation, one can save much more than before given the incomplete Kloosterman sum of the same length, and one can also estimate the incomplete sum non-trivially even though $M$ is smaller than $q^{1/2}$. In particular, this approach works very well when $\varpi$ is around $2/3$. The details can be found in Section 6.5.
@@ 1.4. Connections with Landau–Siegel zeros
We write $\mathrm{e}(t)=\mathrm{e}^{2\pi\mathrm{i}t}$. We use $\varepsilon$ to denote a small positive number, which might be different at each occurrence. The notation $n\sim N$ means $N<n\leqslant 2N.$ We write $A\leqslant B^{-}$ if $A\leqslant B^{1-\delta}$ holds for some constant $\delta>0.$

### v3

@@ Abstract.
The classical Brun–Titchmarsh theorem gives an upper bound, which is of correct order of magnitude in the full range, for the number of primes $p\leqslant x$ satisfying $p\equiv a\,(\operatorname{mod}{q})$. We strength this inequality for $\log q/\log x$ in different ranges, improving upon previous works by Motohashi, Goldfeld, Iwaniec, Friedlander and Iwaniec, and Maynard for general or special moduli. In particular, we are able to beat Iwaniec’s barrier $q<x^{9/20-}$, and improve all the existing inequalities in the range $x^{9/20}\ll q<x^{1/2-}$ by utilizing bilinear or trilinear structures in the remainder terms of linear sieve. The proof is based on various estimates for character and exponential sums, by appealing to arithmetic exponent pairs and bilinear forms with algebraic trace functions from $\ell$-adic cohomology, trilinear forms with Kloosterman fractions, and sums of Kloosterman sums from spectral theory of automorphic forms, as well as large value theorem for Dirichlet polynomials.
@@ 1.1. Main results
In this paper, we establish several choices of $C(\varpi)$ for general moduli and smooth moduli $q\sim x^{\varpi}$. We now quickly state our first result.
@@
Let $\varpi\in[9/20,1/2[.$ Then we may take
(1.4) $C(\varpi)$ $=\frac{66}{33-16\varpi}-C^{*}(\varpi)$
@@
(1.5) $\begin{split}C^{*}(\varpi)&=-\frac{66}{33-16\varpi}\int_{2}^{4}\frac{\log(t-1)}{t}\mathrm{d}t+\frac{8}{4-(3+\vartheta)\varpi}\int_{\frac{8-(7+2\vartheta)\varpi}{4\varpi}}^{\frac{165(4-(3+\vartheta)\varpi)}{4(33-16\varpi)}-\frac{1}{4}}\frac{\log(t-1)}{t}\mathrm{d}t\\ &\ \ \ \ +\frac{8}{4-(1+\vartheta)\varpi}\int_{\max\{\frac{4-(1+\vartheta)\varpi}{2(2-3\varpi)}-\frac{5}{4},~{}2\}}^{\frac{8-(7+2\vartheta)\varpi}{4\varpi}}\frac{\log(t-1)}{t}\mathrm{d}t,\end{split}$
and for prime $q$ we may choose
(1.6) $\begin{split}C^{*}(\varpi)&=-\frac{66}{33-16\varpi}\int_{2}^{4}\frac{\log(t-1)}{t}\mathrm{d}t+\frac{16}{8-7\varpi}\int_{\frac{8-7\varpi}{4\varpi}}^{\frac{165(8-7\varpi)}{8(33-16\varpi)}}\frac{\log(t-1)}{t}\mathrm{d}t\\ &\ \ \ \ +\frac{16}{8-3\varpi}\int_{\max\{\frac{9\varpi}{4(2-3\varpi)},~{}2\}}^{\frac{8-7\varpi}{4\varpi}}\frac{\log(t-1)}{t}\mathrm{d}t.\end{split}$
Here $\vartheta$ denotes the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$ In particular, one may take $\vartheta=7/64$ thanks to Kim and Sarnak [KS03].
Motohashi [Mo73] took $C(\varpi)={16}/(8-3\varpi)$ for $\varpi<1/3,$ and the admissible range was shortly extended by Goldfeld [Go75] to $\varpi<24/71.$ Iwaniec [Iw82] obtained the same constant for all $\varpi<9/20,$ for which he utilized Burgess’s estimate for short character sums:
@@
holds for any non-trivial Dirichlet character modulo $q$ and $r\in\{1,2,3\}$ (see [Bu63, Bu86]). Taking $r=2$, (1.7) implies the so-called Burgess bound $L(\frac{1}{2},\chi)\ll q^{3/16+\varepsilon}$ for Dirichlet $L$-functions. The restriction $\varpi<9/20$ is quite essential in Iwaniec’s approach. On the other hand, the choice $C(\varpi)=4/(2-\varpi)$ for $\varpi\in[9/20,1/2[$ has been obtained implicitly by Iwaniec [Iw82]. This has been illustrated by Baker [Ba96]. One may compare our bounds in Theorem 1.1 with Iwaniec–Baker’s constant. In particular, we may take $C(\frac{9}{20})\approx 2.4814$ for general $q$ and $C(\frac{9}{20})\approx 2.441$ for prime $q$, however Iwaniec–Baker needed to take $C(\frac{9}{20})\approx 2.5806.$
As one may check, the choices (1.5) and (1.6) for $C^{*}(\varpi)$ are always positive for all $\varpi\in[9/20,1/2[$, and Theorem 1.1 improves all previous known bounds even if one takes $C^{*}(\varpi)=0$ in (1.4). The gain of positive choices of $C^{*}(\varpi)$ comes from a suitable Buchstab iteration in the application of sieve methods, so that one may create trilinear structures in the remainder terms with the aid of the well-factorization of the linear sieve weight constructed by Iwaniec [Iw80]. For general moduli, we estimate a twisted moment of character sums (see Lemma 2.6) by appealing to Kloostermania developed by Deshouillers and Iwaniec [DI82]; and in the prime moduli case, the above choice for $C(\varpi)$ is independent from the Ramanujan–Petersson conjecture, and it will follow from another different approach, by appealing to a recent work of Blomer, Humphries, Khan and Milinovich [BHKM20] on twisted fourth moment of Dirichlet $L$-functions to prime moduli (see Lemma 2.7 below).
@@
$C(\varpi)=\begin{cases}8/(5-5\varpi),\quad&\varpi\in[1/2,~{}12/23[,\\ 32/(32-43\varpi),\quad&\varpi\in[1/2,~{}32/61[,\\ 24/(16-17\varpi),&\varpi\in[32/61,~{}8/15[,\\ 48/(40-49\varpi),&\varpi\in[8/15,~{}7/13[,\\ 16/(11-12\varpi),&\varpi\in[7/13,~{}6/11[,\\ 32/(28-35\varpi),&\varpi\in[6/11,~{}4/7[.\end{cases}$
The statementss of Theorems 1.1 and 1.2 are a bit complicated at first glance. The following table presents our choices of $C(\varpi)$, for some specialized $\varpi>1/2,$ and how much we can improve upon Iwaniec [Iw82].
$\varpi$ Our $C(\varpi)$ Iwaniec’s $C(\varpi)$ Improvement
$16/31$ $3.3067$ $3.3514$ $1.4\%$
$12/23$ $3.3455$ $3.4074$ $1.8\%$
$32/61$ $3.3889$ $3.4366$ $1.3\%$
$8/15$ $3.4615$ $3.5294$ $1.9\%$
$7/13$ $3.5254$ $3.5862$ $1.6\%$
$6/11$ $3.5918$ $3.6667$ $2.0\%$
Table 1. Improved $C(\varpi)$
@@
$[9/20,2/3]$ $8/(6-7\varpi)$ Iwaniec [Iw82, Theorem 6]
[9/20,1/2[[9/20,1/2[[ 9 / 20 , 1 / 2 [ $4/(2-\varpi)$ Baker [Ba96, Theorem 1]
@@
Table 2. Admissible values of $C(\varpi)$
The second part of this paper is to reduce the admissible value of $C(\varpi)$ further by assuming that $q$ has good factorizations. For a small constant $\eta>0$, we say $q$ is $\eta$-smooth if all prime factors of $q$ do not exceed $q^{\eta}.$ The smoothness assumption allows us to utilize the factorization of $q$ to control the resultant algebraic exponential sums very effectively. In particular, one may go far beyond the so-called Pólya–Vinogradov barrier for short exponential and character sums, as a consequence of the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs developed by Wu, Xi and Sawin [WX21]. One is referred to Section 3.2 for details.
@@
$C(\varpi)=\frac{24}{15-16\varpi}$
@@
Theorem 1.3 is in fact a special case of the following slightly general formulation on taking $(\kappa,\lambda)=(1/6,2/3)$.
@@
$f(\kappa,\lambda)=\frac{1+\kappa-\lambda}{1+2\kappa-\lambda},\ \ \ g_{\varpi}(\kappa,\lambda)=(3+\kappa-\lambda)-(3+2\kappa-\lambda)\varpi.$
@@
in which case we get $g_{\varpi}(\kappa,\lambda)=0.723659\cdots$. This yields $C(2/3)=4/g_{\varpi}(\kappa,\lambda)\approx 5.52746$ in Theorem 1.4. Note that van Lint and Richert [LR65] obtained the constant $62/(1-\varpi)=$ 6 for $\varpi=2/3.$
Recall the choice of $C(\varpi)$ in Theorem 1.4 and Iwaniec’s choice $8/(6-7\varpi)$ in [Iw82]. We expect that
$\frac{4}{(3+\kappa-\lambda)-(3+2\kappa-\lambda)\varpi}<\frac{8}{6-7\varpi},$
which is equivalent to
$\varpi>\frac{2(\lambda-\kappa)}{1+2\lambda-4\kappa}.$
Taking
$BA^{k-1}B\cdot(0,1)$ $=\Big{(}\frac{1}{2}-\frac{k}{2^{k+1}-2},~{}\frac{1}{2}+\frac{1}{2^{k+1}-2}\Big{)}$
with sufficiently large $k,$ we find Theorem 1.4 improves Iwaniec’s choice $8/(6-7\varpi)$ as long as $\varpi>1/2+\varepsilon$, and $q$ is $\eta$-smooth and squarefree.
The following theorem gives alternative choices for $C(\varpi)$ when $q$ is smooth and $\varpi<1/2.$
@@
Theorem 1.5 can be compared with the following conditional result due to Motohashi [Mo74]. Under the Linderlöf Hypothesis for Dirichlet $L$-functions, he proved that
@@
The following two figures present our new choices of $C(\varpi)$ for $\varpi<1/2$ and $\varpi>1/2,$ respectively, and one may see when we can succeed or not (We choose the very special exponent pair $(\kappa,\lambda)=(1/6,2/3)$ for smooth moduli and $\varpi>1/2$).
@@ 1.2. Ingredients of proof and comparisons with previous works
  • Both of van Lint and Richert [LR65] and Motohashi [Mo74, Mo73] utilized Selberg’s sieve, which is more convenient than Brun’s sieve at that time. Recall that one takes $C(\varpi)=2/(1-\varpi)$ in [LR65], and the constant $2$ is resulted naturally by applications of the linear sieve with its best possible main term. The above denominator $(1-\varpi)$ comes from the level $(x/q)^{1-\varepsilon}$ after a trivial treatment of error terms. It was Motohashi [Mo74, Mo73] who first dealt with the error term in a non-trivial manner, and he successfully employed the bilinear structure in Selberg’s sieve. He employed multiplicative characters, and a basic tool is the mean value of Dirichlet $L$-functions, which prevents him from improving van Lint and Richert [LR65] for $\varpi>1/2.$
  • We also mentioned the works of Montgomery and Vaughan [MV73] and Selberg [Se91] in Table 2. Selberg proved that
    (1.8) $\max_{(a,q)=1}\pi(x;q,a)<\frac{2x}{\varphi(q)(\log(q/x)+c)}$
    holds with $c=2.8$, for all $q<x/c_{0}$ with some $c_{0}>1.$ Selberg employed his own $\Lambda^{2}$ sieve, and published this inequality in [Se91, Page 233]. In another scene, Montgomery and Vaughan [MV73] developed the large sieve, from which they deduced that (1.8) is valid for all $1\leqslant q<x$ and $c=0$.
  • Iwaniec [Iw82] developed a new linear sieve of combinatorial type with well-factorable weights. This far-reaching work was originally motivated by Chen [Ch75] and Motohashi [Mo74], and the merit is that one may optimize the balances between convolutions in the error term, and the sign changes among them can be captured more effectively. In particular, this well-factorization in the error term allows Iwaniec [Iw82] to employ (Weil’s bound for) Kloosterman sums to beat the barrier $\varpi\approx 1/2$. He was able to beat van Lint and Richert [LR65] even if $\varpi$ is close to $2/3-.$ On the other hand, it is worthwhile to mention that Motohashi [Mo99] showed another bilinear structure in the Selberg $\Lambda^{2}$ sieve, which has a similar flavor to that in [Iw82]. In particular, he proved that $C(\varpi)=16/(8-3\varpi)$ is admissible for all $\varpi<9/20,$ which recovered the work of Iwaniec.
@@
In another direction, the coefficient $(c_{r})$ has a convolution structure, coming from Iwaniec’s new construction of the Rosser–Iwaniec sieve. We utilize this special structure of $(c_{r})$ as efficiently as possible, and for $k=1$ we may give another kind of Linderlöf Hypothesis on average for $R\approx q^{(5-2\vartheta)/8}$. Together with large sieves (Lemma 2.1), this can also produce a new level of distribution in linear sieves. The advantage is that $q$ is not necessarily a prime, but we are restricted by the possibility of exceptional spectrum of the non-Eucliean Laplacian. It is also worthwhile to remark that if we may factorize the coefficient $(c_{r})$ as we wish, this argument can produce the same level without appealing to the work of Blomer, Humphries, Khan and Milinovich [BHKM20], modulo the Ramanujan–Petersson conjecture.
@@
(1) If $q$ is squarefree and has only very small prime factors, the factorization of $q$ can be controlled very well, so that the work on arithmetic exponent pairs, developed by Wu, Xi and Sawin [WX21], can be applied. In such situation, one can save much more than before given the incomplete Kloosterman sum of the same length, and one can also estimate the incomplete sum non-trivially even though $M$ is smaller than $q^{1/2}$. In particular, this approach works very well when $\varpi$ is around $2/3$. The details can be found in Section 6.5.
@@ 1.3. Connections with Landau–Siegel zeros
We write $\mathrm{e}(t)=\mathrm{e}^{2\pi\mathrm{i}t}$. We use $\varepsilon$ to denote a small positive number, which might be different at each occurrence. The notation $n\sim N$ means $N<n\leqslant 2N.$ We write $A\leqslant B^{-}$ if $A\leqslant B^{1-\delta}$ holds for some constant $\delta>0.$ Denote by $\varphi,$ $\Lambda$ and $\mu$ the Euler, von Mangoldt and Möbius functions, respectively. For a complex coefficient $\boldsymbol{\alpha}=(\alpha_{n}),$ we define the $\ell_{2}$-norm $\|\boldsymbol{\alpha}\|=(\sum_{n}|\alpha_{n}|^{2})^{\frac{1}{2}}.$